The natural rate of unemployment

Suppose that the markup of the prices of products over wage

cost, z, is 10%, and that the wage-setting equation is

W = P*(1 - 2m + z)

where m is the unemployment rate and z is the unemployment

benefit/minimum wage.

a. What is the real wage, as determined by the price-setting

equation?

b. Solve for the natural rate of unemployment

c. What happens to the natural rate of unemployment if z

falls from 10% to 5%? Explain your answer.
Answer #1

A. Real wage under price setting equation:

W/P=1/(1+z)= 1/(1+0.1)= 0.90

B. Natural rate of unemployment:

W/P=1-2m+z= 0.9

1-2m+0.1=0.9

1-0.8=2m

m=0.1

Natural rate of unemployment=0.1 or 10%

C. If z=0.05

1-2m+0.05=0.9

1-0.05-0.9=2m

0.05/2=m

0.025=m

Natural rate of unemployment =0.025 or 2.5%

When markup falls, natural rate of unemployment falls.